MsBrains (ms-brains):

PLEASE HELP! i will give medal! Simplify and show work so I can learn 7x+4y+3(x-y)-2x Justification {Distributive property, commutative property of addition, combine like terms} Justify your answers!

OpenStudy (alexandervonhumboldt2):

first distribute: 3*(x-y)=???

OpenStudy (alexandervonhumboldt2):

3*(x-y)=3x-3y now we have 7x+4y+3x-3y-2x=combine like terms=y+8x

MsBrains (ms-brains):

This is what the problem is, it's in my study guide

MsBrains (ms-brains):

Nobody wants to help?? Lol